OpenStudy (anonymous):

suppose f(x)= ax^2 + bx + c where a does not equal 0. Show that the vertex of the graph of f is the point (-b/2a, 4ac-b^2/4a)

myininaya (myininaya):

wow I have actually done this one before i wonder if i can find it

myininaya (myininaya):

please hold...

myininaya (myininaya):

well i couldn't find it but thats okay! :)

myininaya (myininaya):

\[y=ax^2+bx+c\] \[ a \neq 0\] So we need to put this in the form \[y=a(x-h)^2+k\] So that means we need to involve completing the square! So we have \[y=ax^2+bx+c\] We need the coefficient of \[x^2\] to be 1. So we can sorta do something like that, but we want the x term with it so we have \[y=a(x^2+\frac{b}{a}x)+c\] How do you feel about this so far?

myininaya (myininaya):

We can complete the square for the thing inside ( )

myininaya (myininaya):

\[y=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a \cdot (\frac{b}{2a})^2\]

myininaya (myininaya):

Whatever you add in you must take out!

myininaya (myininaya):

We added in and took out \[a \cdot (\frac{b}{2a})^2\] In other words we did this exact thing: \[a \cdot (\frac{b}{2a})^2-a \cdot (\frac{b}{2a})^2\]

myininaya (myininaya):

\[y=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a \cdot (\frac{b}{2a})^2 \] \[y=a(x+\frac{b}{2a})^2+c-a \cdot \frac{b^2}{4a^2}\]

myininaya (myininaya):

\[y=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a}\]

myininaya (myininaya):

You can write those two last terms as one fraction

myininaya (myininaya):

You need to find a common denominator!

myininaya (myininaya):

To do that of course.*

myininaya (myininaya):

\[y=a(x+\frac{b}{2a})^2+\frac{4a}{4a} \cdot c-\frac{b^2}{4a}\]

myininaya (myininaya):

I multiply by 1 so now I have the same denominator for the last two terms

myininaya (myininaya):

\[y=a(x+\frac{b}{2a})^2+\frac{4ac-b^2}{4a}\]

myininaya (myininaya):

This is vertex form!

myininaya (myininaya):

Where \[h=\frac{-b}{2a} \text{ and } k=\frac{4ac-b^2}{4a}\] and the vertex is (h,k)